OpenStudy (anonymous):

The tables below show the values of y corresponding to different values of x. Table A x 3 2 1 y 1 0 0 Table B x 3 5 5 y -2 1 2 Which statement is true for the tables?

OpenStudy (anonymous):

Both Table A and Table B represent functions. Both Table A and Table B do not represent functions. Table A does not represent a function but Table B represents a function. Table A represents a function but Table B does not represent a function

OpenStudy (anonymous):

@Purplerainbowcherry @annas @VortexAlliby

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Wouldnt it be D?

OpenStudy (anonymous):

Because table A represents a function, but B doesnt.. Right?

OpenStudy (anonymous):

I was right with D, can you help me on some though?

OpenStudy (anonymous):

a isnt a funtion i dont believe its either B or its or C sorry i mostly meant C but B is possible. i would probably shoot for C

OpenStudy (anonymous):

Alright can you help me with other ones?

OpenStudy (anonymous):

do i get a medal :) O_O

OpenStudy (anonymous):

yes xD

OpenStudy (anonymous):

Which of the following equations best represents the graph of a function f(m)? f(y) = m f(x) = m y = f(m) x = f(m)

OpenStudy (anonymous):

@sloths

OpenStudy (anonymous):

either C or D

OpenStudy (anonymous):

I think its C

OpenStudy (anonymous):

i believe your right

OpenStudy (anonymous):

If f(x) = x + 5, what does f(8) represent? The value of x when (x + 5) = 8 The value of 8f(x) when x = 8 The value of f(x) when (x + 5) = 8 The value of (x + 5) when x = 8

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

Didnt I give you a medal?

OpenStudy (anonymous):

:) :)

OpenStudy (anonymous):

not for that one :)

OpenStudy (anonymous):

how do I medal? lol

OpenStudy (anonymous):

best responce check

OpenStudy (anonymous):

I already did that..

OpenStudy (anonymous):

yayyyyyyyy

OpenStudy (anonymous):

lol.. ok I already posted the third question